• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name rousseau

    Log entry time 08:09:11 on July 31, 2006

    Entry number 176494

    keyword=run summary

    end of run 1914 : Lh2 in, 14microA, momentum 689.2MeV/c, angle 38.78

    radiator moved out helium valve opened on loop2, closed down on loop3. current increased to 20microA run 1915 : LH2 radiatior out target not responding, beam off to reboot run 1915 ended beam back to 20microA run 1916 : LH2, radiator out (again) valve opened on loop3 and closed on loop2 for the changfe of the target target moved to LD2 run 1917 to 1922: LD2, 20microA end of run

    move of the spectrometer to 55.42 it takes something like 2hours, so we'd be glad if someone could have a look at this... change momentum to 0.6421 run 1923 : C optics run 1924 : 4cm dummy run 1925 : 15cm dummy run 1926 : tantalum

    run 1927 : LD2, radiator out, 20microA run 1928 : LH2, radiator out, 20microA put the radiator in, asked for 14microA run 1929 : LH2, radiator in, 14microA shift finished with this run